OpenStudy (anonymous):

find sec[sin^-1(3/5)] in radians

OpenStudy (inkyvoyd):

sec=1/cos sin^2 x+cos^2 x=1

OpenStudy (anonymous):

Ok, but I do not see where that is taking me.

OpenStudy (inkyvoyd):

oh, that's an arcsine isn't it...

OpenStudy (anonymous):

yes.

OpenStudy (inkyvoyd):

this is cheating basically, but wiki lists some arctrig function identities...

OpenStudy (anonymous):

I'll check it out!

OpenStudy (inkyvoyd):

so you're taking the cosine of an arcsine and finding the reciprocal

OpenStudy (anonymous):

Indeed

OpenStudy (inkyvoyd):

=\(\Huge\frac{1}{\sqrt{1-x^2}}\) according to wikipedia

OpenStudy (anonymous):

1/((1-(3/5)^2)^-1)

OpenStudy (anonymous):

^-2 instead of ^-1

OpenStudy (inkyvoyd):

yeh I guess

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

and 5pi/4?

OpenStudy (inkyvoyd):

shrugs, use wolfram alpha to check your work hehe

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks